Determine the currents I_1, I_2, and I_3 in the fi
Answer #1

here,

using KVL in the upper loop

-12 + 1 * I2 + 11 * I2 + 12 * I1 - 12 + ( 22 + 1)* I1 = 0

-12 + 1 * I2 + 11 * I2 + 12 * I1 - 12 + ( 22 + 1)* I1 = 0 ...(1)

using KVL in the lower loop

- 6 + 1 * I3 + 18 * I3 - (11 + 1) * I2 + 12 + 15 * I3 = 0

- 6 + 1 * I3 + 18 * I3 - (11 + 1) * I2 + 12 + 15 * I3 = 0 ...(2)

and

I1 = I2 + I3 ...(3)

from equation(1) , (2) and (3)

I1 = 0.51 A , I2 = 0.51 A

I3 = 0.003 A

B)

the terminal voltage , V = 6 - I3 * r

V = 5.997 V

the terminal voltage is 5.997 V